# Spin pumping into quantum spin nematic states

### v1

@@ Abstract
We theoretically study spin pumping into a spin-nematic state in a junction system composed of a ferromagnetic insulator and a spin-nematic insulator, described by using the spin-1 bilinear-biquadratic model. We analyze an increase of the Gilbert damping in ferromagnetic resonance (FMR) due to an interfacial exchange coupling within a mean-field theory based on the Schwinger boson method. We find that the two Schwinger bosons contribute in distinct ways to spin pumping. We report a detailed dependence of the spin pumping on a resonant frequency, a magnetic field, and an interface type.
@@ I Introduction
In the development of spintronics, the technique of spin pumping driven by ferromagnetic resonance (FMR) has emerged as a versatile tool for injecting spins into materials adjacent to ferromagnets [1, 2, 3, 4, 5, 6]. Spin pumping can also serve as a probe for spin excitations of target materials through the Gilbert damping, which can be extracted from the width of the FMR peak. In fact, enhancement of the Gilbert damping due to interfacial coupling with target materials has been shown to contain valuable information regarding their spin excitations [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27].
@@
Spin pumping into exotic materials with hidden orders represents a significant challenge in identifying novel states of matter and understanding their quantum nature. Among these exotic hidden orders, the spin-nematic state, characterized by spin-rotation symmetry breaking without any accompanying magnetic order, stands as a representative example. The spin-nematic state is described by quadrupolar order or higher multipolar order and has been extensively studied theoretically, using models such as the bilinear-biquadratic exchange model [28, 29, 30, 31, 32, 33, 34, 35, 36, 37] and the $J_{1}$-$J_{2}$ frustrated model [38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51], and the dimer model [52, 53, 54, 55]. These theoretical models predict the existence of ferro- or antiferro-quadrupolar states. The realization of the spin nematic state has also been discussed in a variety of materials such as ${\mathrm{LiCuVO}}_{4}$ [56, 57, 58] and ${\mathrm{BaCdVO}}(\mathrm{PO}4)_{2}$ [59, 60, 61]. Recently, a study of the spin Seebeck effect in the spin-nematic state has also been conducted [62].
@@
In this work, we theoretically calculate the enhancement of Gilbert damping resulting from spin pumping into a spin-nematic insulator. We consider a magnetic junction composed of a spin-nematic (SN) insulator and a ferromagnetic insulator (FI) subjected to microwave irradiation (see Fig. 1). To discuss qualitative features of spin pumping, we employ a simple model based on the bilinear-biquadratic exchange model to realize the spin-nematic state. We derive an analytical expression for the enhancement of the Gilbert damping by second-order perturbation theory with respect to interface coupling and the Schwinger-boson mean-field method. Furthermore, we examine how the enhancement of Gilbert damping depends on the FMR frequency, magnetic field, and interface type.
@@
The rest of this work is organized as follows. In Sec. II, we introduce our theoretical model for a FI/SN junction system and explain the Schwinger-boson mean-field theory for the spin-nematic phase. In Sec. III, we theoretically calculate the enhancement of Gilbert damping due to interfacial exchange coupling within second-order perturbation. In Sec. IV, we show how the increase of Gilbert damping depends on the orientation and strength of the external magnetic field for both clean and dirty interfaces. Finally, we summarize our results in Sec. V. Appendix A is devoted to a detailed derivation of analytic expressions.
@@ II Model
$\mathcal{H}=\mathcal{H}_{\rm FI}+\mathcal{H}_{\rm SN}+\mathcal{H}_{\rm int},$ (1)
where $\mathcal{H}_{\rm FI}$, $\mathcal{H}_{\rm SN}$, and $\mathcal{H}_{\rm int}$ represent FI, SN, and an interfacial exchange coupling. Explicit forms of each Hamiltonian are provided in the following subsections.
@@ II.2 Spin-nematic (SN) state
$\mathcal{H}_{\rm SN}=\sum_{\left<i,j\right>}\left[J_{1}\,{\bm{S}}_{i}\cdot{\bm{S}}_{j}+J_{2}\,({\bm{S}}_{i}\cdot{\bm{S}}_{j})^{2}\right]-h\sum_{i}S^{z^{\prime}}_{i},$ (4)
where ${\bm{S}}_{i}$ denotes a localized spin, $J_{1}$ and $J_{2}$ denote the strengths of the bilinear and biquadratic couplings, respectively, and $h$ represents an external magnetic field. For simplicity in the calculation, we introduced a new coordinate $(x^{\prime},y^{\prime},z^{\prime})$, to the original ordinate as
@@
where ${\bm{r}}_{\delta}$ denotes a displacement vector spanning a site to its neighboring site $\delta$. This Hamiltonian can be diagonalized by the Bogoliubov transformation,
@@
$\varepsilon_{{\bm{k}}}^{\nu}=\frac{zM}{2}\sqrt{A_{{\bm{k}}\nu}^{2}-B_{{\bm{k}}\nu}^{2}},\quad(\nu=z^{\prime},\perp).$ (23)
II.3 Interfacial exchange coupling
We consider the exchange coupling at the FI/SN interface, with Hamiltonian is given as
@@ II.3 Interfacial exchange coupling
where ${\cal S}^{\pm}_{\bm{k}}={\cal S}^{x}_{\bm{k}}\pm i{\cal S}^{y}_{\bm{k}}$ and $S^{\pm}_{\bm{k}}=S^{x}_{\bm{k}}\pm iS^{y}_{\bm{k}}$ are Fourier transformations of the spin ladder operators in the FI and the SN, respectively. Since we focus on the uniform spin precession in the FI, the coupling at the interface is approximated as
@@
We should note that the spin ladder operators are defined in the laboratory coordinates $(x,y,z)$, which are related to the coordinates $(x^{\prime},y^{\prime},z^{\prime})$ as
@@
In our study, we consider two types of the FI/SN interface, i.e., a dirty and clean interface:
${\rm dirty\ interface:}$ $\ T_{{\bm{0}},{\bm{k}}^{\prime}}=\bar{T},$ (27)
${\rm clean\ interface:}$ $\ T_{{\bm{0}},{\bm{k}}^{\prime}}=\bar{T}\delta_{{\bm{k}}^{\prime},{\bm{0}}}.$ (28)
We note that only the uniform spin component in the spin-nematic insulator is excited by spin pumping for the clean interface due to the momentum conservation law.
@@ III Formulation
In this section, we formulate the Gilbert damping in the FMR experiment for the FI/SN bilayer system. First, we briefly describe a general formulation in Sec. III.1. Next, we calculate the increase of the Gilbert damping for the dirty and clean interfaces in Sec. III.2 and Sec. III.3, respectively.
III.1 Gilbert dumping
@@
III.2 Dirty interface
For the model of the dirty interface given in Eq. (27), the self-energy is written as
$\Sigma(\tau)$ $=-\sum_{a,b=1}^{3}\sum_{{\bm{k}},{\bm{k}}^{\prime}}\frac{|\bar{T}|^{2}}{\hbar}g_{a}(\theta)g_{b}(\theta)\left\langle S^{(a)}_{{\bm{k}}}(\tau)\left\{S^{(b)}_{{\bm{k}}^{\prime}}(0)\right\}^{\dagger}\right\rangle.$ (40)
@@ III.2 Dirty interface
$+\cos^{4}(\theta/2)\chi^{+^{\prime}-^{\prime}}_{{\bm{k}},-{\bm{k}}}(\tau)+\sin^{4}(\theta/2)\chi^{-^{\prime}+^{\prime}}_{{\bm{k}},-{\bm{k}}}(\tau)$
$+\frac{1}{4}\sin^{2}\theta\,[\chi^{+^{\prime}+^{\prime}}_{{\bm{k}},-{\bm{k}}}(\tau)+\chi^{-^{\prime}-^{\prime}}_{{\bm{k}},-{\bm{k}}}(\tau)]\Bigr{\}},$ (42)
where we have used translational invariance of ${\cal H}_{\rm SN}$ and $\chi^{z^{\prime}+^{\prime}}=\chi^{z^{\prime}-^{\prime}}=\chi^{+^{\prime}z^{\prime}}=\chi^{-^{\prime}z^{\prime}}=0$. Here, it is remarkable that $\chi^{-^{\prime}-^{\prime}}_{{\bm{k}},-{\bm{k}}}(\tau)$ and $\chi^{+^{\prime}+^{\prime}}_{{\bm{k}},-{\bm{k}}}(\tau)$ does not vanish.
The remaining task is to calculate the spin-spin correlation function $\chi_{{\bm{k}},-{\bm{k}}}^{\mu\nu}(\tau)$ and its Fourier transformation. This is done by simple calculation using the mean-field Hamiltonian, Eq. (22), written in terms of the Schwinger bosons. A detailed calculation is given in Appendix A. Consequently, the increase of the Gilbert damping for $\omega_{\bm{0}}>0$ is given as
@@
+$+\{\cosh 2\xi_{z}(\bm{k})-\cos\mu\sinh 2\xi_{z}(\bm{k})\}\cos^{2}\theta$
$+\frac{1+\cos\mu}{2}\{\cosh 2\xi_{z}(\bm{k})-\sinh 2\xi_{z}(\bm{k})\}\sin^{2}\theta\Bigr{]},$ (44)
@@
III.3 Clean interface
The result for a clean interface is easily obtained by restricting the sum of the wavenumber to ${\bm{k}}={\bm{0}}$. Noting that $\varepsilon_{{\bm{0}}\perp}=0$, the increase of the Gilbert damping for $\omega_{\bm{0}}>0$ is given as
$\delta\alpha_{\rm G}$ $=\frac{2\pi S_{0}|\bar{T}|^{2}}{\hbar\omega_{\bm{0}}}C_{z}^{\prime}\delta(\hbar\omega_{\bm{0}}-\varepsilon_{{\bm{0}}z}),$ (46)
$C_{z}^{\prime}$ $=M\Bigl{[}\sin\mu\cos\theta$
+$+\{\cosh 2\xi_{0}-\cos\mu\sinh 2\xi_{0}\}\cos^{2}\theta$
$+\frac{1+\cos\mu}{2}\{\cosh 2\xi_{0}-\sinh 2\xi_{0}\}\sin^{2}\theta\Bigr{]},$ (47)
where $\xi_{0}=\xi_{{\bm{k}}={\bm{0}}z}$.
IV Result
In this section, we present the results for the increase of the Gilbert damping in the FI/SN bilayer system. First, we show the results in the absence of a magnetic field ($\mu=0$) in Sec. IV.1. Next, we discuss the effect of a finite magnetic field ($\mu>0$) for the dirty and clean interfaces in Sec. IV.2 and Sec. IV.3, respectively. Throughout this paper, the parameters are set as $J_{1}=0$ and $J_{2}=-{\cal J}$, which stabilizes the spin-nematic state. Furthermore, we replace the delta-function in Eqs. (43) and (46) with the Lorentzian function of width $0.03{\cal J}$ to plot the dependence on the FMR frequency $\omega_{\bm{0}}$. We note that $\delta\alpha_{\rm G}$ becomes temperature-independent within the mean-field theory, which is justified at low temperatures.

### v2

@@ Abstract
We theoretically study spin pumping into a spin-nematic state in a junction system composed of a ferromagnetic insulator and a spin-nematic insulator, described by using the spin-1 bilinear-biquadratic model. We analyze the increase of the Gilbert damping in ferromagnetic resonance (FMR) due to an interfacial exchange coupling within a mean-field theory based on the Schwinger boson method. We find that the two Schwinger bosons contribute in distinct ways to spin pumping. We report a detailed dependence of the spin pumping on a resonant frequency and a magnetic field, along with a comparison to the case of the canted antiferromagnetic insulator. Our findings establish spin pumping as a relevant technique to verify and characterize spin-nematic phases.
@@ I Introduction
In the development of spintronics, the technique of spin pumping driven by ferromagnetic resonance (FMR) has emerged as a versatile tool for injecting spins into materials adjacent to ferromagnets [1, 2, 3, 4, 5, 6]. Spin pumping can also serve as a probe for spin excitations of target materials through the Gilbert damping, which can be extracted from the width of the FMR peak. In fact, enhancement of the Gilbert damping due to interfacial coupling with target materials has been shown to contain valuable information regarding their spin excitations [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32].
@@
Spin pumping into exotic materials with hidden orders represents a significant challenge in identifying novel states of matter and understanding their quantum nature. Among these exotic hidden orders, the spin-nematic state, characterized by spin-rotation symmetry breaking without any accompanying magnetic order, stands as a representative example. The spin-nematic state is described by quadrupolar order or higher multipolar order and has been extensively studied theoretically, using models such as the bilinear-biquadratic exchange model [33, 34, 35, 36, 37, 38, 39, 40, 41, 42], the $J_{1}$-$J_{2}$ frustrated model [43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56], and the dimer model [57, 58, 59, 60]. These theoretical models predict the existence of ferro- or antiferro-quadrupolar states. The realization of the spin nematic state has also been discussed in a variety of materials such as ${\mathrm{LiCuVO}}_{4}$ [61, 62, 63] and ${\mathrm{BaCdVO}}(\mathrm{PO}4)_{2}$ [64, 65, 66]. In the absence of dipole moments, the unique properties of the SN state are difficult to observe using static magnetization measurements. Conventional approaches employ NMR relaxation rates in an attempt to identify SN by focusing on a feature appearing in the low-temperature behavior. However, this feature is challenging to detect experimentally, due to the difficulty in separating the two magnetic excitations [67]. Recently, a study of the spin Seebeck effect in the spin-nematic state has also been conducted [68].
@@
In this work, we theoretically calculate the enhancement of the Gilbert damping resulting from spin pumping into a spin-nematic insulator. We consider a magnetic junction composed of a spin-nematic (SN) insulator and a ferromagnetic insulator (FI) subjected to microwave irradiation (see Fig. 1). To discuss qualitative features of spin pumping, we employ a simple model based on the bilinear-biquadratic exchange model to realize the spin-nematic state. We derive an analytical expression for the enhancement of the Gilbert damping by the Schwinger-boson mean-field method and second-order perturbation theory with respect to interface coupling. Furthermore, we examine how the enhancement of Gilbert damping depends on the FMR frequency and magnetic field. For comparison with other magnetic states, we also calculate the enhancement of the Gilbert damping resulting from spin pumping into a canted antiferromagnetic insulator, and discuss the difference to the spin nematic.
@@
The rest of this work is organized as follows. In Sec. II, we introduce our theoretical model for junction systems and explain the Schwinger-boson mean-field theory for the spin-nematic phase. In Sec. III, we theoretically calculate the enhancement of the Gilbert damping due to interfacial exchange coupling within second-order perturbation. In Sec. IV, we show how the increase of the Gilbert damping depends on the orientation and strength of the external magnetic field and compare its results with that of a canted antiferromagnetic insulator. Finally, we summarize our results in Sec. V. Appendix A and appendix B are devoted to a detailed derivation of analytic expressions.
@@ II Model
$\mathcal{H}_{{\rm FI}/{\rm SN}}$ $=\mathcal{H}_{\rm FI}+\mathcal{H}_{\rm SN}+\mathcal{H}_{\rm int},$ (1)
where $\mathcal{H}_{\rm FI}$, $\mathcal{H}_{\rm SN}$, and $\mathcal{H}_{\rm int}$ represent FI, SN, and an interfacial exchange coupling, respectively. Explicit forms of each Hamiltonian are provided in the following subsections.
To identify the feature of the nematic phase through the FMR experiment, we also consider a junction composed of FI and a canted antiferromagnetic insulator (CAF). We note that the SN and CAF phases have a few common features; (i) they have two excitation modes and (ii) they do not have the U(1) symmetry, leading to the breakdown of the conservation law of a magnon number (or a spin component in a specific direction). The model Hamiltonian of the FI/CAF junction is expressed as
$\mathcal{H_{{\rm FI}/{\rm CAF}}}$ $=\mathcal{H}_{\rm FI}+\mathcal{H}_{\rm CAF}+\mathcal{H}_{\rm int},$ (2)
where $H_{\rm cant}$ describes CAF and $\mathcal{H}_{\rm int}$ represents an interfacial exchange coupling with the same form as the FI/SN junction system. By comparing the results for the SN/FI and CAF/FI junctions, we clarify how to identify the feature of the SN phase in the FMR experiment.
@@ II.2 Spin-nematic (SN) state
[…]
$\hskip 14.22636pt-h\sum_{i}S^{z^{\prime}}_{i}-D\sum_{i}(S^{x^{\prime}}_{i})^{2},$ (5)
where ${\bm{S}}_{i}$ denotes a localized spin, $J_{1}$ and $J_{2}$ denote the strengths of the bilinear and biquadratic couplings, respectively, and $h$ represents an external magnetic field. For simplicity in the calculation, we assume $D$ is much smaller than $J_{1}$, $J_{2}$, and $h$, existing solely to determine the direction of quadrupolar ordering. We employ a triangular lattice following a previous study [39] though spin nematic states can appear also in a square lattice. We introduced a new coordinate $(x^{\prime},y^{\prime},z^{\prime})$, to the original ordinate as
@@
where ${\bm{r}}_{\delta}$ denotes a displacement vector spanning a site to its neighboring site $\delta$. Note the formal similarity of the terms with $B_{{\bm{k}}\nu}$ to quantum squeezing known in quantum optics which has been pointed out by  [7]. This Hamiltonian can be diagonalized by the Bogoliubov transformation,
@@
$\varepsilon_{{\bm{k}}\nu}=\frac{zM}{2}\sqrt{A_{{\bm{k}}\nu}^{2}-B_{{\bm{k}}\nu}^{2}},\quad(\nu=z^{\prime},\perp).$ (24)
II.3 Canted antiferromagnetic insulator (CAF)
Figure 2: Coordinate transformation between the coordinates $O$-$x^{\prime}y^{\prime}z^{\prime}$ and the canted coordinates for each sublattices (a) $O$-$x_{\rm A}y_{\rm A}z_{\rm A}$ and (b) $O$-$x_{\rm B}y_{\rm B}z_{\rm B}$.
To describe the CAF insulator, we consider the three-dimensional Heisenberg model on a cubic lattice, whose Hamiltonian is expressed as
$\mathcal{H}_{\rm CAF}=\sum_{\left<i,j\right>}J_{3}\,{\bm{S}}_{i}\cdot{\bm{S}}_{j}-h\sum_{i}S^{z^{\prime}}_{i}-D\sum_{i}(S^{x^{\prime}}_{i})^{2},$ (25)
where ${\bm{S}}_{i}$ denotes a localized spin, $J_{3}$ ($>0$) is an antiferromagnetic exchange coupling, $h$ represents an external magnetic field, and the coordinate $(x^{\prime},y^{\prime},z^{\prime})$ is defined by Eq. (6) (see also Fig. 1(b)). We divide the lattice into A and B sublattices and further assume that $D$ is much smaller than $J_{3}$ and $h$, existing solely to determine the direction of magnetic ordering. The spin at each site points to the $+z^{\prime}$ direction for a strong magnetic field while it points to the ±$\pm x^{\prime}$ direction for a weak magnetic field due to the small anisotropy $D$, depending on the sublattice. To interpolate the rotation of the ordered spin as a function of the magnetic field, we introduce an angle $\phi$ and new coordinates, $(x_{\rm A},y_{\rm A},z_{\rm A})$ and $(x_{\rm B},y_{\rm B},z_{\rm B})$ as shown in Fig. 2(a) and (b):
$\begin{pmatrix}x^{\prime}\\ y^{\prime}\\ z^{\prime}\end{pmatrix}=\begin{pmatrix}\cos\phi&0&-\sin\phi\\ 0&1&0\\ \sin\phi&0&\cos\phi\end{pmatrix}\begin{pmatrix}x_{\rm A}\\ y_{\rm A}\\ z_{\rm A}\end{pmatrix},$ (26)
$\begin{pmatrix}x^{\prime}\\ y^{\prime}\\ z^{\prime}\end{pmatrix}=\begin{pmatrix}\cos\phi&0&\sin\phi\\ 0&1&0\\ -\sin\phi&0&\cos\phi\end{pmatrix}\begin{pmatrix}x_{\rm B}\\ y_{\rm B}\\ z_{\rm B}\end{pmatrix}.$ (27)
In these coordinates, the averaged spin is represented as $(S_{\rm A}^{x},S_{\rm A}^{y},S_{\rm A}^{z})=(0,0,S)$ and $(S_{\rm B}^{x},S_{\rm B}^{y},S_{\rm B}^{z})=(0,0,S)$, where $S$ is an amplitude of the localized spin ${\bm{S}}_{i}$. Using the Holstein-Primakoff transformation, we introduce magnons for each sublattice as follows,
$S^{+}_{{\rm A},i}$ $=\sqrt{2S}a_{i},\ S^{-}_{{\rm A},i}=\sqrt{2S}a^{\dagger}_{i},\ S^{z}_{{\rm A},i}=S-a^{\dagger}_{i}a_{i}$ (28)
$S^{+}_{{\rm B},i}$ $=\sqrt{2S}b_{i},\ S^{-}_{{\rm B},i}=\sqrt{2S}b^{\dagger}_{i},\ S^{z}_{{\rm B},i}=S-b^{\dagger}_{i}b_{i}.$ (29)
With the Fourier transformations of the magnon operators, $a_{\bm{k}}$ and $b_{\bm{k}}$, we introduce new magnon operators as follows:
$\begin{pmatrix}a_{\bm{k}}\\ b_{-\bm{k}}\end{pmatrix}=\frac{1}{\sqrt{2}}\begin{pmatrix}1&1\\ -1&1\end{pmatrix}\begin{pmatrix}c_{\bm{k}}\\ d_{-\bm{k}}\end{pmatrix}.$ (30)
Leaving the leading-order terms of $1/S$, the Hamiltonian is approximately rewritten as
${\cal H}_{\rm CAF}$ $=\sum_{\bm{k}}\Bigl{\{}A_{\bm{k}}c_{{\bm{k}}}^{\dagger}c_{{\bm{k}}}-\frac{C_{\bm{k}}}{2}(c_{{\bm{k}}}c_{{-\bm{k}}}+c_{{\bm{k}}}^{\dagger}c^{\dagger}_{{-\bm{k}}})$
$\hskip 14.22636pt+B_{\bm{k}}d_{{\bm{k}}}^{\dagger}d_{{\bm{k}}}+\frac{C_{\bm{k}}}{2}(d_{{\bm{k}}}d_{{-\bm{k}}}+d_{{\bm{k}}}^{\dagger}d^{\dagger}_{{-\bm{k}}})\Bigr{\}}.$ (31)
Here, the coefficients are given as
$A_{\bm{k}}$ $=6J_{3}S(1-2\cos^{2}\phi)+h\cos\phi$
$\hskip 14.22636pt+2J_{3}S\gamma_{\bm{k}}(1-\cos^{2}\phi),$ (32)
$B_{\bm{k}}$ $=6J_{3}S(1-2\cos^{2}\phi)+h\cos\phi$
$\hskip 14.22636pt-2J_{3}S\gamma_{\bm{k}}(1-\cos^{2}\phi),$ (33)
$C_{\bm{k}}$ $=2J_{3}S\gamma_{\bm{k}}\cos^{2}\phi,$ (34)
where $\gamma_{\bm{k}}=\cos k_{x}+\cos k_{y}+\cos k_{z}$. As for the SN state the terms with $C_{\bm{k}}$ lead to squeezing of the magnons. To diagonalize the Hamiltonian, we consider the following Bogoliubov transformation
$\begin{pmatrix}c_{\bm{k}}\\ c^{{\dagger}}_{-\bm{k}}\end{pmatrix}=\begin{pmatrix}\cosh\psi_{\bm{k}}^{\alpha}&\sinh\psi_{\bm{k}}^{\alpha}\\ \sinh\psi_{\bm{k}}^{\alpha}&\cosh\psi_{\bm{k}}^{\alpha}\end{pmatrix}\begin{pmatrix}\alpha_{\bm{k}}\\ \alpha^{{\dagger}}_{-\bm{k}}\end{pmatrix},$ (35)
$\begin{pmatrix}d_{\bm{k}}\\ d^{{\dagger}}_{-\bm{k}}\end{pmatrix}=\begin{pmatrix}\cosh\psi_{\bm{k}}^{\beta}&\sinh\psi_{\bm{k}}^{\beta}\\ \sinh\psi_{\bm{k}}^{\beta}&\cosh\psi_{\bm{k}}^{\beta}\end{pmatrix}\begin{pmatrix}\beta_{\bm{k}}\\ \beta^{{\dagger}}_{-\bm{k}}\end{pmatrix},$ (36)
where $\psi^{\nu}_{{\bm{k}}}$ is chosen to satisfy
$\tanh(2\psi^{\alpha}_{\bm{k}})$ $=\frac{C_{\bm{k}}}{A_{\bm{k}}},$ (37)
$\tanh(2\psi^{\beta}_{\bm{k}})$ $=\frac{-C_{\bm{k}}}{B_{\bm{k}}}.$ (38)
$\mathcal{H}_{\rm CAF}=\sum_{\bm{k}}\varepsilon_{{\bm{k}}\alpha}\alpha_{\bm{k}}^{\dagger}\alpha_{\bm{k}}+\sum_{\bm{k}}\varepsilon_{{\bm{k}}\beta}\beta_{\bm{k}}^{\dagger}\beta_{\bm{k}},$ (39)
where $\varepsilon_{\nu}(\bm{k})$ is a dispersion of the two bosonic modes given as
$\varepsilon_{{\bm{k}}\alpha}$ $=\Big{[}2J_{3}S(3-\gamma_{\bm{k}})\{2J_{3}S(3+\gamma_{\bm{k}})$
$\hskip 19.91692pt+h\cos\phi-4J_{3}S(3+\gamma_{\bm{k}})\cos^{2}\phi\}\Big{]}^{1/2},$ (40)
$\varepsilon_{{\bm{k}}\beta}$ $=\Big{[}2J_{3}S(3+\gamma_{\bm{k}})\{2J_{3}S(1-2\cos^{2}\phi)(3-\gamma_{\bm{k}})$
$\hskip 19.91692pt+h\cos\phi\}\Big{]}^{1/2}.$ (41)
II.4 Interfacial exchange coupling
We consider the exchange coupling at the interface, with Hamiltonian is given as
@@ II.4 Interfacial exchange coupling
where ${\cal S}^{\pm}_{\bm{k}}={\cal S}^{x}_{\bm{k}}\pm i{\cal S}^{y}_{\bm{k}}$ and $S^{\pm}_{\bm{k}}=S^{x}_{\bm{k}}\pm iS^{y}_{\bm{k}}$ are Fourier transformations of the spin ladder operators in the FI and the SN or the canted antiferromagnetic insulator, respectively. Since we focus on the uniform spin precession in the FI, the coupling at the interface is approximated as
@@
For simplicity, we set $T_{{\bm{0}},{\bm{k}}^{\prime}}=\bar{T}$, assuming a dirty interface [70, 15, 71, 72]. We should note that the spin ladder operators are defined in the laboratory coordinates $(x,y,z)$, which are related to the coordinates $(x^{\prime},y^{\prime},z^{\prime})$ as
@@ III Formulation
In this section, we formulate the Gilbert damping in the FMR experiment for the bilayer system. First, we briefly describe a general formulation in Sec. III.1. Next, we calculate the increase of the Gilbert damping into the SN and CAF insulators in Sec. III.2 and Sec. III.3, respectively.
III.1 Gilbert damping
@@
III.2 Spin pumping into SN insulator
For the model of the SN state, the self-energy is written as
$\Sigma(\tau)$ $=-\sum_{a,b=1}^{3}\sum_{{\bm{k}},{\bm{k}}^{\prime}}\frac{|\bar{T}|^{2}}{\hbar}g_{a}(\theta)g_{b}^{*}(\theta)\left\langle S^{(a)}_{{\bm{k}}}(\tau)\left\{S^{(b)}_{{\bm{k}}^{\prime}}(0)\right\}^{\dagger}\right\rangle.$ (56)
@@ III.2 Spin pumping into SN insulator
$+\cos^{4}\left(\frac{\theta}{2}\right)\chi^{+^{\prime}-^{\prime}}_{{\bm{k}},-{\bm{k}}}(\tau)+\sin^{4}\left(\frac{\theta}{2}\right)\chi^{-^{\prime}+^{\prime}}_{{\bm{k}},-{\bm{k}}}(\tau)$
$+\frac{\sin^{2}\theta}{4}[\chi^{+^{\prime}+^{\prime}}_{{\bm{k}},-{\bm{k}}}(\tau)+\chi^{-^{\prime}-^{\prime}}_{{\bm{k}},-{\bm{k}}}(\tau)]\Bigr{\}},$ (58)
where we have used translational invariance of ${\cal H}_{\rm SN}$ and $\chi^{z^{\prime}+^{\prime}}_{{\bm{k}},{\bm{k}}^{\prime}}=\chi^{z^{\prime}-^{\prime}}_{{\bm{k}},{\bm{k}}^{\prime}}=\chi^{+^{\prime}z^{\prime}}_{{\bm{k}},{\bm{k}}^{\prime}}=\chi^{-^{\prime}z^{\prime}}_{{\bm{k}},{\bm{k}}^{\prime}}=0$. Here, it is remarkable that $\chi^{-^{\prime}-^{\prime}}_{{\bm{k}},-{\bm{k}}}(\tau)$ and $\chi^{+^{\prime}+^{\prime}}_{{\bm{k}},-{\bm{k}}}(\tau)$ do not vanish as a consequence of the squeezing terms in Eq. (14).
The remaining task is to calculate the spin-spin correlation function $\chi_{{\bm{k}},-{\bm{k}}}^{\mu\nu}(\tau)$ and its Fourier transformation. This is done by using the mean-field Hamiltonian, Eq. (23), written in terms of the Schwinger bosons. The detailed calculation is given in Appendix A. Consequently, the increase of the Gilbert damping for $\omega_{\bm{0}}>0$ is given as
@@
+$+\{\cosh 2\xi_{{\bm{k}}z}-\cos\mu\sinh 2\xi_{{\bm{k}}z}\}\cos^{2}\theta$
$+\frac{1+\cos\mu}{2}\{\cosh 2\xi_{{\bm{k}}z}-\sinh 2\xi_{{\bm{k}}z}\}\sin^{2}\theta\Bigr{]},$ (60)
@@
III.3 Spin pumping into CAF insulator
Similarly, the self-energy is calculated for the model of the CAF insulator as
$\Sigma(\tau)$ $=\sum_{{\bm{k}}}\sum_{{\bm{k}^{\prime}}=\pm{\bm{k}}}\frac{|\bar{T}|^{2}}{\hbar}\Bigl{\{}\sin^{2}\theta\,\chi^{z^{\prime}z^{\prime}}_{{\bm{k}},{\bm{k}^{\prime}}}(\tau)$
$+\cos^{4}\left(\frac{\theta}{2}\right)\chi^{+^{\prime}-^{\prime}}_{{\bm{k}},{\bm{k}^{\prime}}}(\tau)+\sin^{4}\left(\frac{\theta}{2}\right)\chi^{-^{\prime}+^{\prime}}_{{\bm{k}},{\bm{k}^{\prime}}}(\tau)$
$+\frac{\sin^{2}\theta}{4}[\chi^{+^{\prime}+^{\prime}}_{{\bm{k}},{\bm{k}^{\prime}}}(\tau)+\chi^{-^{\prime}-^{\prime}}_{{\bm{k}},{\bm{k}^{\prime}}}(\tau)]\Bigr{\}},$ (62)
where $\chi^{\mu\nu}_{{\bm{k}},{\bm{k}^{\prime}}}(\tau)$ ($\mu,\nu=z^{\prime},+^{\prime},-^{\prime}$) is an imaginary-time spin-spin correlation function
$\chi^{\mu\nu}_{{\bm{k}},{\bm{k}^{\prime}}}(\tau)=-\frac{1}{\hbar}\sum_{\eta={\rm A},{\rm B}}\left\langle S^{\mu}_{{\eta,\bm{k}}}(\tau)S^{\nu}_{{\eta,\bm{k}}^{\prime}}(0)\right\rangle.$ (63)
We have used translational invariance of ${\cal H}_{\rm CAF}$ and $\chi^{z^{\prime}+^{\prime}}_{{\bm{k}},{\bm{k}}^{\prime}}=\chi^{z^{\prime}-^{\prime}}_{{\bm{k}},{\bm{k}}^{\prime}}=\chi^{+^{\prime}z^{\prime}}_{{\bm{k}},{\bm{k}}^{\prime}}=\chi^{-^{\prime}z^{\prime}}_{{\bm{k}},{\bm{k}}^{\prime}}=0$. We note that $\chi^{-^{\prime}-^{\prime}}_{{\bm{k}},-{\bm{k}}}(\tau)$ and $\chi^{+^{\prime}+^{\prime}}_{{\bm{k}},-{\bm{k}}}(\tau)$ do not vanish as in the SN insulator, reflecting the absence of the U(1) symmetry. The spin-spin correlation function $\chi_{{\bm{k}},{\bm{k}^{\prime}}}^{\mu\nu}(\tau)$ is obtained from the diagonalized Hamiltonian, Eq. (41). Consequently, the increase of the Gilbert damping for $\omega_{\bm{0}}>0$ is given as
$\delta\alpha_{\rm G}$ $=\frac{2\pi S_{0}|\bar{T}|^{2}}{\hbar\omega_{\bm{0}}}\sum_{\bm{k}}\sum_{\nu=\alpha,\beta}C_{\nu}\delta(\hbar\omega_{\bm{0}}-\varepsilon_{{\bm{k}}\nu}),$ (64)
$C_{\nu}$ $=M\Bigl{[}\cos\phi\cos\theta$
$+\frac{1}{2}\left\{(\cos^{2}\phi+\sin^{2}\phi\sin^{2}\theta)(\cosh 2\psi_{\bm{k}}^{\nu}+\sinh 2\psi_{\bm{k}}^{\nu})\right\}$
$+\frac{1}{2}\left\{(1-\sin^{2}\theta)(\cosh 2\psi_{\bm{k}}^{\nu}-\sinh 2\psi_{\bm{k}}^{\nu})\right\}\Bigr{]}.$ (65)
A detailed calculation is given in Appendix B.
IV Results
In this section, we present the results for the increase of the Gilbert damping in the FI/SN and FI/CAF bilayer systems. First, we show the results in the absence of a magnetic field ($h=0$) and under a finite magnetic field ($h>0$) in Sec. IV.1 and Sec. IV.2, respectively. Throughout this paper, the parameters are set as $J_{1}=0$, $J_{2}=-{\cal J}$, and $J_{3}={\cal J}$, where this choice of $J_{1}$ and $J_{2}$ realizes the spin nematic phase. Furthermore, we replace the delta-function in Eqs. (59) and (64) with the Lorentzian function of width $0.03{\cal J}$ to plot the dependence on the FMR frequency $\omega_{\bm{0}}$. We note that $\delta\alpha_{\rm G}$ becomes temperature-independent within the mean-field theory, which is justified at low temperatures.